If you make an initial contribution of $8000 to a retirement account 5 years from now and make annual contributions of $8000 thereafter, after how many contributions will your balance have reached at least $1000000 if your account earns 6% annual interest?
Answer #1

Here we can use the Future Value of Annuity (FVA) formula:

|(1 + i)n-1 FVA = A

Where A = Annuity / Contribution

i = rate of interest

n = number of contributions

So let us put the values in the formula,

(1 +0.06) -1 1,000,000 = 8000 0.06

125=(1.06 -1 0.06

7.5 = (1.06) - 1

8.5 = (1.06)

8.5 = 1.06

n = 36

It doesn't make difference whether you contribute from today or 5 years later, number of contributions will remain same.

This 36 contributions include the initial contribution.

Verification:

Year Opening Interest @ 6% on opening Contribution Balance 1,528 2,100 8,000 8,000 8,000 8,000 8,000 8,000 8,000 2,706 3,348

Year Opening Contribution Balance Interest @ 6% on opening 12,323 13,542 14,835 16,205 17,657 19,197 20,828 22,558 8,000 8,00

If you contribute $8000 from 5 years from now at the rate of 6% p.a, then your account will show balance of $1,010,145 at the end of 41st year. Total contributions including initial contribution is 36.